# Modeling temporal hypergraphs

### v1

@@ Abstract
Networks representing social, biological, technological or other systems are often characterized by higher-order interaction involving any number of nodes. Temporal hypergraphs are given by ordered sequences of hyperedges representing sets of nodes interacting at given points in time. In this paper we discuss how a recently proposed model family for time-stamped hyperedges – relational hyperevent models (RHEM) – can be employed to define tailored null distributions for temporal hypergraphs. RHEM can be specified with a given vector of temporal hyperedge statistics – functions that quantify the structural position of hyperedges in the history of previous hyperedges – and equate expected values of these statistics with their empirically observed values. This allows, for instance, to analyze the overrepresentation or underrepresentation of temporal hyperedge configurations in a model that reproduces the observed distributions of possibly complex sub-configurations, including but going beyond node degrees. Concrete examples include, but are not limited to, preferential attachment, repetition of subsets of any given size, triadic closure, homophily, and degree assortativity for subsets of any order.
I Introduction
@@
Graphs – collections of nodes and edges, where each edge connects two nodes – can represent networks of interconnected entities [1, 2]. Hypergraphs [3] generalize graphs by containing hyperedges that can connect any number of nodes. While having been suggested for network analysis decades ago [4], hypergraphs received a recent surge of interest [5, 6, 7, 8, 9, 10], in part due the insight that crucial information gets lost when forcing hyperedges into collections of dyadic edges [11, 12]. Examples of complex systems that can be represented by hypergraphs are numerous and include networks of meetings or social events, emails, coauthoring and citation, co-offending, and chemical reactions [13, 14, 15, 16, 17, 18, 19]. Many of these examples give rise to temporal networks [20, 21], in this case, hypergraphs whose hyperedges are time-stamped or time-ordered [22].
@@
Quantitative properties of real-world networks are often compared to expected quantities in random graphs that reproduce some observed properties but randomize others. For example, if a network has more closed triangles and a skewed degree distribution compared to uniform random graphs [23, 24], then it is informative to analyze whether the degree distribution alone can already explain the higher number of triangles, or whether the latter comprises an additional significant observation beyond what is implied by node degrees. For this purpose one might employ random graphs that reproduce the observed degree sequence (exactly or in expectation) but that randomize everything else [25] – a type of random graph models that can also be defined for random hypergraphs with given hyperedge sizes and node degrees (bipartite configuration model). Random models, for graphs or for hypergraphs, [25, 5, 26, 11, 27] may serve different purposes, which however are interrelated. Frequent use cases employ random graph models (i) as null distributions to compare observed values with expected values, e. g., [28], (ii) to find or test network effects, such as preferential attachment, homophily, reciprocity, assortative mixing, or triadic closure, e. g., [29, 30, 31], (iii) to generate random (hyper)graphs, e. g., [32, 33, 34, 27], or (iv) to predict (hyper)graphs, e. g., [35, 32, 36, 37].
@@
In this paper we review a recently proposed family of random models for temporal hypergraphs – relational hyperevent models (RHEM) [12, 38] – and illustrate their use to define tailored null distributions and to test network effects in temporal hypergraphs. RHEM can be specified with a given vector of temporal hyperedge statistics – functions that quantify how hyperedges are positioned in the history of previous hyperedges – and successively model the probability of the next hyperedge to reproduce in expectation the observed statistics. We argue that RHEM are a valuable addition to available random models for temporal hypergraphs (see a comprehensive list in [8]) due to the following properties. (i) RHEM are a versatile model family that can be specified with any given vector of temporal hypergraph pattern, including but not limited to preferential attachment, partial repetition of subsets of any order, homophily, assortative mixing, or triadic closure. This allows, for instance, to analyze the overrepresentation or underrepresentation of temporal hyperedge configurations in a model that reproduces the observed distributions of potentially complex sub-configurations. (ii) RHEM can be fitted to real-world networks to equate expected values and observed values in the given vector of temporal hyperedge statistics, as formalized in Eq. (4). (iii) RHEM can be generalized in several directions, including directed hypergraphs, multilayer or multimode hypergraphs, and to hypergraphs with typed, labeled, signed, or weighted hyperedges; compare Sect. IV. (iv) From a statistical perspective, RHEM belong to the family of Cox regression models [39], which gives access to a rich set of theoretical results and available software [40, 41, 42].
@@
II Preliminaries
II.1 Hypergraphs and temporal hypergraphs
A hypergraph is a pair $(\mathcal{I},H)$, where =$\mathcal{I}=\{i_{1},\dots,i_{N}\}$ is a set of nodes (also denoted as actors in our main illustrating example) and $H\subseteq\mathcal{P}(\mathcal{I})$ is a set of hyperedges, that is, subsets of the node set. The size of a hyperedge $I\in H$ is the number of its nodes (i. e., $size(I)=|I|$) and the degree of a node $i\in\mathcal{I}$ in a hypergraph $(\mathcal{I},H)$ is the number of hyperedges containing $i$ (i. e., $deg(i)=|\{I\in H\colon i\in I\}|$). More generally, the degree of a set of nodes $I\subseteq\mathcal{I}$ is the number of hyperedges containing all nodes in $I$ (i. e., $deg(I)=|\{I^{\prime}\in H\colon I\subseteq I^{\prime}\}|$).
@@ II.1 Hypergraphs and temporal hypergraphs
$E=[e_{1}=(t_{1},I_{1}),\dots,e_{M}=(t_{M},I_{M})]\enspace,$
where, for each $m=1,\dots,M$, it is $I_{m}\subseteq\mathcal{I}$. We note that “time” may, but does not have to, represent “time shown on a clock”. For instance, time may also just be a counter marking the order of arrival of hyperedges. Throughout most of this paper, we assume that time stamps are strictly ordered ($t_{1}<t_{2}<\dots<t_{M}$), that is, there are no simultaneous events. (Note, however, the discussion on “tied event times” given in Sect. IV.) We further note that hyperedges may be exactly repeated in $E$, that is, it may be $I_{m}=I_{m^{\prime}}$ for different indices $m\neq m^{\prime}$.
@@
$deg(I,E_{<t})=|\{(t_{m},I_{m})\in E_{<t}\colon I\subseteq I_{m}\}|\enspace.$
@@
II.2 Data: Les Misérables
For illustration we use as a running example the network of actor co-appearance in chapters of Victor Hugo’s novel Les Misérables, compiled by Donald Knuth [51]. The data comprises 80 actors (co-)appearing in one or several of 288 chapters. Each of these chapters defines one hyperevent $e=(t,I)$, whose hyperedge $I$ contains the actors starring in that chapter and where the time stamps are counters for chapters $t\in\{1,\dots,288\}$ in the order in which they appear in the novel. To illustrate the use of exogenous node attributes, we augment Knuth’s Les Misérables data by a binary gender variable, $female(i)$, which is one if and only if actor $i$ is female. We believe that, using actors’ names, Knuth’s short description, and our own reading of the novel, we can derive this variable with good reliability. (We note that our analysis is given for illustrating the model and not for making any claims about the demography of the novel’s characters.) Among the 80 actors, there are 28 female (35%). The Les Misérables data is convenient for illustrating models in this paper, since it is established in the networks literature, e. g., [53, 54], and since the actors and their relations are part of a story, allowing to anecdotally interpret certain findings. We list applications of RHEM to several other real-world temporal hypergraphs in Sect. IV.
III RHEM for temporal hypergraphs
III.1 Model framework
@@
$\mathbb{E}_{M_{0}}\left[degsum(\cdot,E_{<t})\right]=|I|\sum_{I_{m}\in E_{<t}}\frac{|I_{m}|}{|\mathcal{I}|}\enspace.$
If, however, we assume a preferential attachment mechanism in which nodes with higher prior degrees are selected into the next hyperedge with higher probability, then the observed value of $degsum$ would likely be higher than the expected value under the uniform null model, providing empirical evidence for preferential attachment of individual nodes. The RHEM family, defined below, can be used to fit a model $M_{1}$ such that the expected value of $degsum$ over the entire sequence of events under $M_{1}$ is equal to the observed value and this model, in turn, can be used to test whether preferential attachment of individual nodes is already sufficient to explain other patterns in temporal hypergraphs, such as preferential attachment of higher-order sets, assortativity, or triadic closure. In the following we review the definition of RHEM and illustrate its use as a general model family for temporal hypergraphs designed to test hypothetical patterns, simultaneously controlling for a given vector of temporal hyperedge statistics $\bm{x}(\cdot,E_{<t_{m}})$.
@@
$P(I_{m}|E_{<t_{m}},\bm{\beta})=\frac{\exp[\bm{\beta}\cdot\bm{x}(I_{m},E_{<t_{m}})]}{\sum_{I\in\binom{\mathcal{I}}{|I_{m}|}}\exp[\bm{\beta}\cdot\bm{x}(I,E_{<t_{m}})]}\enspace.$ (3)
@@
Conveniently, the specification in Eqs. (2) and (3) is identical to common specifications of the partial likelihood of Cox proportional hazard (CoxPH) models [39, 40, 38], one of the most established models for event history analysis. This gives the exponential of the parameters, $\exp(\beta_{h})$, the interpretation as “hazard ratios”, that is, factors by which the event rate on a set of nodes $I$ is expected to increase (or decrease), if the associated hyperedge statistic $x_{h}(I,E_{<t})$ increases by one unit. The connection to the CoxPH model also implies an abundance of existing software packages to estimate parameters and a wealth of theoretical results. Of special importance for RHEM are results about the consistency of parameters estimated from sampled likelihoods – obtained by case-control sampling [56, 38] –, where the index set $\binom{\mathcal{I}}{|I_{m}|}$ in the denominator of (3) gets replaced by a sampled set containing the hyperedge $I_{m}$ of the observed event (the “case”) plus $K$ non-event hyperedges (“controls”) $\{I_{1},\dots,I_{K}\}$ which are sampled uniformly at random from $\binom{\mathcal{I}}{|I_{m}|}$. Estimation via sampled likelihoods is necessary since the entire set $\binom{\mathcal{I}}{|I_{m}|}$ is often prohibitively large for all but the smallest node sets $\mathcal{I}$ and hyperedge sizes $|I_{m}|$.
Exponential-family models, like the one specified in Eqs. (2) and (3), have the property that the observed explanatory variables are equal to their expectation in the model with parameters estimated by maximizing the likelihood. (This well-know fact can be derived by setting the partial derivatives of the log-likelihood to zero.) Concretely, for the maximum likelihood estimates $\hat{\bm{\beta}}$ it is
$\sum_{m=1}^{M}\bm{x}(I_{m},E_{<t_{m}})=\sum_{m=1}^{M}\mathbb{E}_{\hat{\bm{\beta}}}\left[\bm{x}(\cdot,E_{<t_{m}})\right]\enspace.$ (4)
@@
III.2 Commonly used temporal hyperedge statistics
@@
$subrep^{(k)}(I,E_{<t})=\sum_{I^{\prime}\in\binom{I}{k}}deg(I^{\prime},E_{<t})\enspace.$ (5)
(The index set $\binom{I}{k}$ is the set of all subsets of size $k$, that is, $\binom{I}{k}=\{I^{\prime}\subseteq I\colon|I^{\prime}|=k\}$.) We point out that the previously discussed $degsum$ statistic is identical to $subrep^{(1)}$. Figure 1 illustrates subset repetition of order one, two, and three. For example, subset repetition of order one takes higher values on the hyperedge $I^{\prime}=\{D,E,F\}$ while subset repetition of order three is higher for $I=\{A,B,C\}$, while both hyperedges have the identical value for subset repetition of order two.
@@
Figure 1: (Adapted from [12].) Illustrating subset repetition based on four prior events $e_{1},\dots,e_{4}$ on the hyperedges $I=\{A,B,C\}$ and $I^{\prime}=\{D,E,F\}$ at time $t>t_{4}$. For $I$ subset repetition of order one, two, and three is equal to one. For $I^{\prime}$ it is $subrep^{(1)}(I^{\prime},E_{<t})=2$, $subrep^{(2)}(I^{\prime},E_{<t})=1$, and $subrep^{(3)}(I^{\prime},E_{<t})=0$.
@@
closure $0.008\;(0.003)^{**}$ $0.006\;(0.003)^{*}$
assort(1) $0.010\;(0.002)^{***}$ $0.010\;(0.002)^{***}$
assort(2) $0.006\;(0.002)^{***}$ $0.005\;(0.002)^{**}$
female $-0.118\;(0.112)$
@@
We can also standardize differences between observed and expected values, dividing by the estimated standard deviations, to obtain a z-score measure. For example, we get a z-score of 5.4 for the JV-CO dyad and a z-score of 6.7 for MA-CO, so that both dyads are significantly overrepresented. On the other hand, the z-score of JV-MA is $-0.45$, so that this dyad is not significantly underrepresented but the difference to the expected value could be explained by random chance.
Turning to overrepresentation or underrepresentation of filled triads, according to Model 1, we find that the most overrepresented triad is among three of the “friends of the ABC”, namely EN-CM-CR, which co-appear in 12 chapters (the maximum value over all triads) and are expected to co-appear in only 0.01 chapters. The triad formed by the three most frequent actors JV-MA-CO is expected to appear in 3.3 chapters but does appear in 6 chapters in the actual novel, placing this triad on position 164 (among 82,160 triads) when ranking triads by their overrepresentation according to Model 1. The assessment of the JV-MA-CO triad changes remarkably if we additionally control for a preferential attachment effect of actor pairs. According to Model 2 in Table 3, JV-MA-CO is the most underrepresented triad, with an expected number of 16 co-appearances but only 6 co-appearances in the actually observed data. This difference to the expectation of Model 1 can be explained by the finding that two of the three contained dyads, namely JV-CO and MA-CO, are strongly overrepresented (controlling for individual frequencies). Considering also previous dyadic co-appearances, the probability that all three, JV, MA, and CO, jointly appear in a chapter increases. This illustrates that if we want to assess the relative frequency of higher-order structures (such as filled triads), we should not only control for the prior degree of individual nodes but also for the prior occurrence of more complex sub-structures, such as dyads that are subsets of the respective triads. We can find a similar reasoning in [52] who argue that to test higher-order homophily in hypergraphs one has to control for lower-order homophily.
A variation of subset repetition is exact repetition, which captures a tendency to exactly repeat the participant list of prior events. This is a frequent pattern, for example, in email networks [38] (where senders repeatedly send messages to the same list of receivers) or in temporal hypergraphs representing meetings [43] (where regular meetings may have fixed attendee lists). We can test this effect with the statistic $repetition(I,E_{<t})$ counting the number of prior events on the hyperedge $I$
$repetition(I,E_{<t})=|\{(t_{m},I_{m})\in E_{<t}\colon I=I_{m}\}|\enspace.$ (6)
@@
$\sum_{\{i_{1},i_{2}\}\in\binom{I}{2}\atop i_{3}\in\mathcal{I}\colon i_{3}\neq i_{1},i_{2}}\min[deg(\{i_{1},i_{3}\}),deg(\{i_{2},i_{3}\})]\enspace,$ (7)
where we abbreviate $deg(I^{\prime},E_{<t})$ by $deg(I^{\prime})$ For each dyad $\{i_{1},i_{2}\}$ in $I$ we look for “third nodes” $i_{3}$ (not necessarily in $I$) and then we weight the two-path $i_{1}$$i_{3}$$i_{2}$ by the minimum prior degree of $\{i_{1},i_{3}\}$ and $\{i_{2},i_{3}\}$. We then add up these weights over all dyads included in $I$ and all third nodes. Models in the second row of Table 3 (also see Models 7 and 9 in Table 1) reveal that closure events are overrepresented and that this finding is robust if we control for repetition of individual nodes, dyads, and triads.
@@
Degree assortativity is the pattern that nodes tend to connect with other nodes having a similar degree. As it is the case for preferential attachment (implemented in RHEM via subset repetition), assortativity in temporal hypergraphs can by defined for subsets of any size, not just for individual nodes. We define a temporal hyperedge statistic assortativity of order $k$, for varying $k\geq 1$, by
@@
where we abbreviate $deg(I^{\prime},E_{<t})$ by $d(I^{\prime})$. The summation index runs over all unordered pairs $I^{\prime},I^{\prime\prime}$ of different subsets of size $k$ of the given hyperedge $I$. For each of these pairs we add the absolute difference between the degrees $d(I^{\prime})$ and $d(I^{\prime\prime})$. The resulting sum would be a measure of disassortativity as it gets larger when the prior degrees of the contained subsets are more different. We therefore take the negative sum to get a measure of degree assortativity. It is not necessary to add a constant in order to make the statistic positive, since any additive constant would cancel out in (3). We note that $assort^{(k)}(I,E_{<t})$ can assume non-zero values only if $|I|\geq k+1$, since there have to be more than one subset of size $k$ to get any pairs of such subsets.

### v2

@@ Abstract
Networks representing social, biological, technological or other systems are often characterized by higher-order interaction involving any number of nodes. Temporal hypergraphs are given by ordered sequences of hyperedges representing sets of nodes interacting at given points in time. In this paper we discuss how a recently proposed model family for time-stamped hyperedges – relational hyperevent models (RHEM) – can be employed to define tailored null distributions for temporal hypergraphs and to test and control for complex dependencies in hypergraph dynamics. RHEM can be specified with a given vector of temporal hyperedge statistics – functions that quantify the structural position of hyperedges in the history of previous hyperedges – and equate expected values of these statistics with their empirically observed values. This allows, for instance, to analyze the overrepresentation or underrepresentation of temporal hyperedge configurations in a model that reproduces the observed distributions of possibly complex sub-configurations, including but going beyond node degrees. Concrete examples include, but are not limited to, preferential attachment, repetition of subsets of any given size, triadic closure, homophily, and degree assortativity for subsets of any order.
Keywords: higher-order networks, temporal hypergraphs, random graph models, relational event models
1Department of Computer and Information Science, University of Konstanz, Germany
2Department of Sociology, University of Bucharest, Panduri 90-92, Bucharest, 050663, Romania
3The Center for Innovation in Medicine, Theodor Pallady Blv. 42J, Bucharest, 032266, Romania
4Faculty of Natural Sciences and Mathematics, University of Maribor, Koroška cesta 160, 2000 Maribor, Slovenia
5Community Healthcare Center Dr. Adolf Drolc Maribor, Ulica talcev 9, 2000 Maribor, Slovenia
6Department of Physics, Kyung Hee University, 26 Kyungheedae-ro, Seoul 02447, Republic of Korea
7University College, Korea University, 145 Anam-ro, Seoul 02841, Republic of Korea
1 Introduction
@@
Graphs – collections of nodes and edges, where each edge connects two nodes – can represent networks of interconnected entities [1, 2]. Hypergraphs [3] generalize graphs by containing hyperedges that can connect any number of nodes. While having been suggested for network analysis decades ago [4], hypergraphs received a recent surge of interest [5, 6, 7, 8, 9, 10], in part due to the insight that crucial information gets lost when forcing hyperedges into collections of dyadic edges [11, 12]. Examples of complex systems that can be represented by hypergraphs are numerous and include networks of meetings or social events, emails, coauthoring and citation, co-offending, and chemical reactions [13, 14, 15, 16, 17, 18, 19]. Many of these examples give rise to temporal networks [20, 21], in this case, hypergraphs whose hyperedges are time-stamped or time-ordered [22].
@@
Quantitative properties of real-world networks are often compared to expected quantities in random graphs that reproduce some observed properties but randomize others. For example, if a network has more closed triangles and a skewed degree distribution compared to uniform random graphs [23, 24], then it is informative to analyze whether the degree distribution alone can already explain the higher number of triangles, or whether the latter comprises an additional significant observation beyond what is implied by node degrees. For this purpose one might employ random graphs that reproduce the observed degree sequence (exactly or in expectation) but that randomize everything else [25] – a type of random graph models that can also be defined for random hypergraphs with given hyperedge sizes and node degrees (bipartite configuration model). Random models, for graphs or for hypergraphs, [25, 5, 26, 11, 27, 28] may serve different purposes, which however are interrelated. Frequent use cases employ random graph models (i) as null distributions to compare observed values with expected values, e. g., [29], (ii) to find or test network effects, such as preferential attachment, homophily, reciprocity, assortative mixing, or triadic closure, e. g., [30, 31, 32], (iii) to generate random (hyper)graphs, e. g., [33, 34, 35, 27], or (iv) to predict (hyper)graphs, e. g., [36, 33, 37, 38].
@@
In this paper we review a recently proposed family of random models for temporal hypergraphs – relational hyperevent models (RHEM) [12, 39] – and illustrate their use to define tailored null distributions and to test network effects in temporal hypergraphs. RHEM can be specified with a given vector of temporal hyperedge statistics – functions that quantify how hyperedges are positioned in the history of previous hyperedges – and successively model the probability of the next hyperedge to reproduce the observed statistics in expectation. We argue that RHEM are a valuable addition to available random models for temporal hypergraphs (see a comprehensive list in [8]) due to the following properties. (i) RHEM are a versatile model family that can be specified with any given vector of temporal hypergraph pattern, including but not limited to preferential attachment, partial repetition of subsets of any order, homophily, assortative mixing, or triadic closure. This allows, for instance, to analyze the overrepresentation or underrepresentation of temporal hyperedge configurations in a model that reproduces the observed distributions of potentially complex sub-configurations. (ii) RHEM can be fitted to real-world networks to equate expected values and observed values in the given vector of temporal hyperedge statistics, as formalized in Eq. (4). (iii) RHEM can be generalized in several directions, including directed hypergraphs, multilayer or multimode hypergraphs, and to hypergraphs with typed, labeled, signed, or weighted hyperedges; compare Sect. 4. (iv) From a statistical perspective, RHEM belong to the family of Cox regression models [40], which gives access to a rich set of theoretical results and available software [41, 42, 43].
@@
2 Preliminaries
2.1 Hypergraphs and temporal hypergraphs
A hypergraph is a pair $(\mathcal{I},H)$, where $\mathcal{I}=\{i_{1},\dots,i_{N}\}$ is a set of nodes (also denoted as actors in our main illustrating example) and $H\subseteq\mathcal{P}(\mathcal{I})$ is a set of hyperedges, that is, subsets of the node set. The size of a hyperedge $I\in H$ is the number of its nodes (i. e., $size(I)=|I|$) and the degree of a node $i\in\mathcal{I}$ in a hypergraph $(\mathcal{I},H)$ is the number of hyperedges containing $i$ (i. e., $deg(i)=|\{I\in H\colon i\in I\}|$). More generally, the degree of a set of nodes $I\subseteq\mathcal{I}$ is the number of hyperedges containing all nodes in $I$ (i. e., $deg(I)=|\{I^{\prime}\in H\colon I\subseteq I^{\prime}\}|$).
@@ 2.1 Hypergraphs and temporal hypergraphs
$E=[e_{1}=(t_{1},I_{1}),\dots,e_{M}=(t_{M},I_{M})]\kern 5.0pt,$
where, for each $m=1,\dots,M$, it is $I_{m}\subseteq\mathcal{I}$. We note that “time” may, but does not have to, represent “time shown on a clock”. For instance, time may also just be a counter marking the order of arrival of hyperedges. Throughout most of this paper, we assume that time stamps are strictly ordered ($t_{1}<t_{2}<\dots<t_{M}$), that is, there are no simultaneous events. (Note, however, the discussion on “tied event times” given in Sect. 4.) We further note that hyperedges may be exactly repeated in $E$, that is, it may be $I_{m}=I_{m^{\prime}}$ for different indices $m\neq m^{\prime}$.
@@
$deg(I,E_{<t})=|\{(t_{m},I_{m})\in E_{<t}\colon I\subseteq I_{m}\}|\kern 5.0pt.$
@@
2.2 Data: Les Misérables
For illustration we use as a running example the network of actor co-appearance in chapters of Victor Hugo’s novel Les Misérables, compiled by Donald Knuth [52]. The data comprises 80 actors (co-)appearing in one or several of 288 chapters. Each of these chapters defines one hyperevent $e=(t,I)$, whose hyperedge $I$ contains the actors starring in that chapter and where the time stamps are counters for chapters $t\in\{1,\dots,288\}$ in the order in which they appear in the novel. To illustrate the use of exogenous node attributes, we augment Knuth’s Les Misérables data by a binary gender variable, $female(i)$, which is one if and only if actor $i$ is female. We believe that, using actors’ names, Knuth’s short description, and our own reading of the novel, we can derive this variable with good reliability. (We note that our analysis is given for illustrating the model and not for making any claims about the demography of the novel’s characters.) Among the 80 actors, there are 28 female (35%). For reproducibility, we provide the preprocessed data, augmented by the femalefemale variable, in the file jean_events.csv, linked from the eventnet Webpage  [^1] https://github.com/juergenlerner/eventnet/tree/master/data/les_miserables.
The Les Misérables data is convenient for illustrating models in this paper, since it is established in the networks literature, e. g., [54, 55], and since the actors and their relations are part of a story, allowing to anecdotally interpret certain findings. We list applications of RHEM to several other real-world temporal hypergraphs in Sect. 4.
3 RHEM for temporal hypergraphs
3.1 Model framework
@@
$\mathbb{E}_{M_{0}}\left[degsum(\cdot,E_{<t})\right]=|I|\sum_{I_{m}\in E_{<t}}\frac{|I_{m}|}{|\mathcal{I}|}\kern 5.0pt.$
If, however, we assume a preferential attachment mechanism in which nodes with higher prior degrees are selected into the next hyperedge with higher probability, then the observed value of degsumdegsum would likely be higher than the expected value under the uniform null model, providing empirical evidence for preferential attachment of individual nodes. The RHEM family, defined below, can be used to fit a model $M_{1}$ such that the expected value of degsumdegsum over the entire sequence of events under $M_{1}$ is equal to the observed value and this model, in turn, can be used to test whether preferential attachment of individual nodes is already sufficient to explain other patterns in temporal hypergraphs, such as preferential attachment of higher-order sets, assortativity, or triadic closure. In the following we review the definition of RHEM and illustrate its use as a general model family for temporal hypergraphs designed to test hypothetical patterns, simultaneously controlling for a given vector of temporal hyperedge statistics $\bm{x}(\cdot,E_{<t_{m}})$.
@@
$P(I_{m}|E_{<t_{m}},\bm{\beta})=\frac{\exp[\bm{\beta}\cdot\bm{x}(I_{m},E_{<t_{m}})]}{\sum_{I\in\binom{\mathcal{I}}{|I_{m}|}}\exp[\bm{\beta}\cdot\bm{x}(I,E_{<t_{m}})]}\kern 5.0pt.$ (3)
@@
Conveniently, the specification in Eqs. (2) and (3) is identical to common specifications of the partial likelihood of Cox proportional hazard (CoxPH) models [40, 41, 39], one of the most established models for event history analysis. This gives the exponential of the parameters, $\exp(\beta_{h})$, the interpretation as “hazard ratios”, that is, factors by which the event rate on a set of nodes $I$ is expected to increase (or decrease), if the associated hyperedge statistic $x_{h}(I,E_{<t})$ increases by one unit. The connection to the CoxPH model also implies an abundance of existing software packages to estimate parameters and a wealth of theoretical results. Of special importance for RHEM are results about the consistency of parameters estimated from sampled likelihoods – obtained by case-control sampling [57, 39] –, where the index set $\binom{\mathcal{I}}{|I_{m}|}$ in the denominator of (3) gets replaced by a sampled set containing the hyperedge $I_{m}$ of the observed event (the “case”) plus $K$ non-event hyperedges (“controls”) $\{I_{1},\dots,I_{K}\}$ which are sampled uniformly at random from $\binom{\mathcal{I}}{|I_{m}|}$. Estimation via sampled likelihoods is necessary since the entire set $\binom{\mathcal{I}}{|I_{m}|}$ is often prohibitively large for all but the smallest node sets $\mathcal{I}$ and hyperedge sizes $|I_{m}|$. Using case-control sampling, RHEM have been applied to networks containing more than a million of nodes and events, e. g., [46].
Exponential-family models, like the one specified in Eqs. (2) and (3), have the property that the observed explanatory variables are equal to their expectation in the model with parameters estimated by maximizing the likelihood. (This well-known fact can be derived by setting the partial derivatives of the log-likelihood to zero.) Concretely, for the maximum likelihood estimates $\hat{\bm{\beta}}$ it is
$\sum_{m=1}^{M}\bm{x}(I_{m},E_{<t_{m}})=\sum_{m=1}^{M}\mathbb{E}_{\hat{\bm{\beta}}}\left[\bm{x}(\cdot,E_{<t_{m}})\right]\kern 5.0pt.$ (4)
@@
3.2 Commonly used temporal hyperedge statistics
@@
$subrep^{(k)}(I,E_{<t})=\sum_{I^{\prime}\in\binom{I}{k}}deg(I^{\prime},E_{<t})\kern 5.0pt.$ (5)
(The index set $\binom{I}{k}$ is the set of all subsets of size $k$, that is, $\binom{I}{k}=\{I^{\prime}\subseteq I\colon|I^{\prime}|=k\}$.) We point out that the previously discussed degsumdegsum statistic is identical to $subrep^{(1)}$. Figure 1 illustrates subset repetition of order one, two, and three. For example, subset repetition of order one takes higher values on the hyperedge $I^{\prime}=\{D,E,F\}$ while subset repetition of order three is higher for $I=\{A,B,C\}$, while both hyperedges have the identical value for subset repetition of order two.
@@
Figure 1: (Adapted from [12].) Illustrating subset repetition based on four prior events $e_{1},\dots,e_{4}$ on the hyperedges $I=\{A,B,C\}$ and $I^{\prime}=\{D,E,F\}$ at time $t>t_{4}$. For II it is $subrep^{(1)}(I^{\prime},E_{<t})=3$, $subrep^{(2)}(I^{\prime},E_{<t})=3$, and $subrep^{(3)}(I^{\prime},E_{<t})=1$. For $I^{\prime}$ it is $subrep^{(1)}(I^{\prime},E_{<t})=6$, $subrep^{(2)}(I^{\prime},E_{<t})=3$, and $subrep^{(3)}(I^{\prime},E_{<t})=1$.
@@
[…]
assort(1) $^{***} 0.010(0.002)0.010\;(0.002)^{***}$
[…]
female 0.118(0.112)-0.118\;(0.112)
@@
We can also standardize differences between observed and expected values, dividing by the estimated standard deviations, to obtain a z-score measure. For example, we get a z-score of 5.4 for the JV-CO dyad and a z-score of 6.7 for MA-CO, so that both dyads are significantly overrepresented. On the other hand, the z-score of JV-MA is 0.45-0.45, so that this dyad is not significantly underrepresented but the difference to the expected value could be explained by random chance.
Turning to overrepresentation or underrepresentation of filled triads, according to Model 1, we find that the most overrepresented triad is among three of the “friends of the ABC”, namely EN-CM-CR, which co-appear in 12 chapters (the maximum value over all triads) and are expected to co-appear in only 0.01 chapters. The triad formed by the three most frequent actors JV-MA-CO is expected to appear in 3.3 chapters but does appear in 6 chapters in the actual novel, placing this triad on position 164 (among 82,160 triads) when ranking triads by their overrepresentation according to Model 1. The assessment of the JV-MA-CO triad changes remarkably if we additionally control for a preferential attachment effect of actor pairs. According to Model 2 in Table 3 (Appendix), specified with subset repetition of order one and two, JV-MA-CO is the most underrepresented triad, with an expected number of 16 co-appearances but only 6 co-appearances in the actually observed data. This difference to the expectation of Model 1 can be explained by the finding that two of the three contained dyads, namely JV-CO and MA-CO, are strongly overrepresented (controlling for individual frequencies). Considering also previous dyadic co-appearances, the probability that all three, JV, MA, and CO, jointly appear in a chapter increases. This illustrates that if we want to assess the relative frequency of higher-order structures (such as filled triads), we should not only control for the prior degree of individual nodes but also for the prior occurrence of more complex sub-structures, such as dyads that are subsets of the respective triads. We can find a similar reasoning in [53] who argue that to test higher-order homophily in hypergraphs one has to control for lower-order homophily.
A variation of subset repetition is exact repetition, which captures a tendency to exactly repeat the participant list of prior events. This is a frequent pattern, for example, in email networks [39] (where senders repeatedly send messages to the same list of receivers) or in temporal hypergraphs representing meetings [44] (where regular meetings may have fixed attendee lists). We can test this effect with the statistic $repetition(I,E_{<t})$ counting the number of prior events on the hyperedge II
$repetition(I,E_{<t})=|\{(t_{m},I_{m})\in E_{<t}\colon I=I_{m}\}|\kern 5.0pt.$ (6)
@@
$\sum_{\{i_{1},i_{2}\}\in\binom{I}{2}\atop i_{3}\in\mathcal{I}\colon i_{3}\neq i_{1},i_{2}}\min[deg(\{i_{1},i_{3}\},E_{<t}),deg(\{i_{2},i_{3}\},E_{<t})]\kern 5.0pt.$ (7)
For each dyad $\{i_{1},i_{2}\}$ in $I$ we look for “third nodes” $i_{3}$ (not necessarily in $I$) and then we weight the two-path $i_{1}$$i_{3}$$i_{2}$ by the minimum prior degree of $\{i_{1},i_{3}\}$ and $\{i_{2},i_{3}\}$. We then add up these weights over all dyads included in $I$ and all third nodes. Models in the second row of Table 3 in the appendix (also see Models 7 and 9 in Table 1) reveal that closure events are overrepresented and that this finding is robust if we control for repetition of individual nodes, dyads, and triads.
@@
Degree assortativity is the pattern that nodes tend to connect with other nodes having a similar degree. As it is the case for preferential attachment (implemented in RHEM via subset repetition), assortativity in temporal hypergraphs can be defined for subsets of any size, not just for individual nodes. We define a temporal hyperedge statistic assortativity of order $k$, for varying $k\geq 1$, by
@@
The summation index runs over all unordered pairs $I^{\prime},I^{\prime\prime}$ of different subsets of size $k$ of the given hyperedge $I$. For each of these pairs we add the absolute difference between the degrees $deg(I^{\prime},E_{<t})$ and $deg(I^{\prime\prime},E_{<t})$. The resulting sum would be a measure of disassortativity as it gets larger when the prior degrees of the contained subsets are more different. We therefore take the negative sum to get a measure of degree assortativity. It is not necessary to add a constant in order to make the statistic positive, since any additive constant would cancel out in (3). We note that $assort^{(k)}(I,E_{<t})$ can assume non-zero values only if $|I|\geq k+1$, since there have to be more than one subset of size $k$ to get any pairs of such subsets.